OpenStudy (anonymous):

4rs-2s-3(rs+1)-2s Simplify.

OpenStudy (anonymous):

distribute -3 to get 4rs -2s-3r+3-2s

OpenStudy (anonymous):

put matching together 4rs-3rs-2s-2s-3

OpenStudy (anonymous):

combine like terms 1rs -4s+4 of rs-4s+4

OpenStudy (anonymous):

I got 1rs+(-3) as my answer

OpenStudy (anonymous):

you made an "operations with integers" mistake show me your steps

OpenStudy (anonymous):

4rs-2s-3(rs+1)-2s I distributed 3 to rs & 1. 4rs-2s-3rs+(-3)-2s I combined like terms (4rs+(-3rs)= 1rs & -2s-2s= 0 So my answer is 1rs+(-3)

OpenStudy (anonymous):

-2s-2s does not equal 0

OpenStudy (anonymous):

but I did make mistakes too here its is

OpenStudy (anonymous):

4rs -2s-3rs+3-2s so (4rs-3rs) = 1rs (-2s-2s)=-4s 3 =3 so its rs-4s+3 that is a common mistake that students thing -2s-2s = 0 because they remember to negatives make a positive but that is only when you multiply them

OpenStudy (anonymous):

can you help me on my next one?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

8y-2z-(y-z)+3yz

OpenStudy (anonymous):

8y-2z-(y-z)+3yz when you see -(y-z) it means -1(y-z) so you distrib the -1

OpenStudy (anonymous):

so to get 8y-2z-y+z+3yz

OpenStudy (anonymous):

now combine like terms